# Less for More: Enhanced Feedback-aligned Mixed LLMs for Molecule Caption Generation and Fine-Grained NLI Evaluation

### v2

@@ Abstract
Scientific language models drive research innovation but require extensive fine-tuning on large datasets. This work enhances such models by improving their inference and evaluation capabilities with minimal or no additional training. Focusing on molecule caption generation, we explore synergies between alignment fine-tuning and model merging in a cross-modal setup. We reveal intriguing insights into the behaviour and suitability of such methods while significantly surpassing state-of-the-art models. Moreover, we propose a novel atomic-level evaluation method leveraging off-the-shelf Natural Language Inference (NLI) models for use in the unseen chemical domain. Our experiments demonstrate that our evaluation operates at the right level of granularity, effectively handling multiple content units and subsentence reasoning, while widely adopted NLI methods consistently misalign with assessment criteria.
@@
Dimitris Gkoumas1,2, Maria Liakata2,3 1Huawei Noah’s Ark Lab, London, UK, 2Queen Mary University of London, London, UK, 3The Alan Turing Institute, London, UK Correspondence: d.gkoumas@qmul.ac.uk
@@ 1 Introduction
AI in Chemistry is essential for developing scalable and cost-effective scientific solutions, such as pioneering drugs Ferguson and Gray (2018), advanced materials Kippelen and Brédas (2009), and improved chemical processes Zhong et al. (2023). The vast search spaces in which these solutions reside make chemical language models crucial for accelerating scientific discovery AI4Science and Quantum (2023); Zhang et al. (2023). Recent trends have led to the use of multimodal models to learn molecular and linguistic representations, either in separate but coordinated spaces Edwards et al. (2021, 2022); Liu et al. (2023a), in a common space Liu et al. (2023b), or through dual approaches Luo et al. (2023); Christofidellis et al. (2023). These models often rely heavily on extensive supervised fine-tuning. However, merely increasing model size and data does not guarantee improvement Tirumala et al. (2022); Xu et al. (2023). Thus we propose focusing on novel training methods to improve molecule language models.
@@
Here we enhance molecule language models using minimal training by leveraging synergies between alignment fine-tuning Ouyang et al. (2022) and model merging Yang et al. (2024) in a crossmodal setup. Specifically, we focus on molecule-language translation, using as little as 10% of the training data Edwards et al. (2024). Figure 1 illustrates our proposed three-step process.
Figure 1: A diagram illustrating the three-step process to enhance molecule language models with minimal training. Black arrows indicate that no training data is required to operate the models, while blue arrows show that only a minimal amount of training data (10%) is used.
@@
Model merging, a technique for fusing models fine-tuned on different tasks, builds a versatile model without needing the original training data or expensive computation. This method has been quickly adopted in foundation models and Large Language Models (LLMs) Yang et al. (2024). We extend this concept to a crossmodal setting by merging per-task pretrained molecule language models (step 1 in Fig. 1), deploying both weight- and subspace-based techniques to obtain universal models (§ 3.2.1).
@@
For fine-tuning alignment, we focus on Reinforcement Learning from Human Feedback (RLHF)Stiennon et al. (2020) to align the universal models. Although alignment has typically been used to calibrate LLM behaviour Askell et al. (2021), we hypothesise that it can also accelerate learning in crossmodal spaces by rewarding preferred over dispreferred outputs, thus improving inference with minimal training data. We focus on optimisation algorithms using closed-form losses on offline preferences, such as Direct Preference Optimisation (DPO) Rafailov et al. (2024), Contrastive Preference Optimisation (CPO) Xu et al. (2024), and Kahneman-Tversky Optimisation (KTO) Ethayarajh et al. (2024). We incorporate golden data as human preferences and dispreferred synthetic outputs generated by proprietary models into the reward signal (Step2 in Fig. 1).
@@
We evaluate our models on out-of-distribution data using established statistical-based metrics Sets (2022); Edwards et al. (2022). Additionally, we use Natural Language Inference (NLI) models to assess generated text within the chemical domain. However, we argue that off-the-shelf NLI models are suboptimal for several reasons: a) they are trained on relatively short texts Williams et al. (2018), while generated text may aggregate multiple content units that partially overlap with different sentences in the reference text Nenkova et al. (2007); b) they are limited by the data they were trained on, making them unreliable for unseen domains McIntosh et al. (2024); and c) they lack subsentence inference, hindering their ability to handle reordered content in generated text (see Fig. 3). Thus we propose a novel atomic-level cross-NLI approach that addresses these issues. By decomposing reference and generated texts into atomic premises and hypotheses using an LLM, we calculate probability distributions of contradiction and entailment via an NLI model and finally apply row-wise operations to obtain novel hallucination and coverage metrics (§$\S$3.3). Our findings and contributions are as follows:
@@ 2.1 LLMs for Chemistry
Existing approaches for LLMs in the chemical domain typically depend on costly pretraining with large unimodal datasets for reaction prediction and retrosynthesis Schwaller et al. (2019); Vaucher et al. (2020), or task-specific fine-tuning for language-molecule learning Edwards et al. (2021, 2022, 2024) and molecule editing Liu et al. (2023a); Fang et al. (2023). Other initiatives focus on multitask learning, which involves resource-intensive pretraining and large multitask datasets Lu and Zhang (2022); Ross et al. (2022); Christofidellis et al. (2023); Zhang et al. (2024). By contrast, we investigate synergies between fine-tuning alignment and model merging to enhance molecule language models with minimal training.
@@ 2.2 Model Merging
Existing model merging techniques can be broadly categorised into weight-based, subspace-based, and routing-based approaches. Advanced weight-based methods typically utilise optimisation algorithms Yang et al. (2023); Akiba et al. (2024) or geometric interpolations Zhou et al. (2024); Goddard et al. (2024) to determine optimal task vector coefficients while preserving the capabilities of independent models. Subspace-based methods involve pruning Yadav et al. (2023); Yu et al. (2024) or masking Wang et al. (2024) to eliminate insignificant parameters before merging sparse models in the parameter subspace, thereby reducing task interference. By contrast, routing-based methods adaptively combine models during inference, adjusting the merge dynamically per the specific input Muqeeth et al. (2023); Tang et al. (2024). We experiment with weight- and subspace-based merging in a crossmodal context.
@@ 2.3 Aligning LLMs
LLM alignment methods can be divided into two categories: test-time and fine-tuning. Test-time alignment techniques, such as prompt engineering and guided decoding Khanov et al. (2024); Huang et al. (2024), adjust LLMs without altering their weights, although performance relies on the original model. Fine-tuning methods, such as RLHF Stiennon et al. (2020); Ouyang et al. (2022), are effective but complex, requiring model training and continuous sampling. DPO Rafailov et al. (2024) simplifies RLHF by focusing on a direct optimisation objective from PPO Schulman et al. (2017), thereby streamlining the process to the training of supervised policy models. CPO Xu et al. (2024) enhances efficiency by utilising a uniform reference model, which reduces memory demands. Other methods employ SFT to optimise the management of the RLHF model and parameter tuning Ethayarajh et al. (2024); Meng et al. (2024). Here, we explore alignment fine-tuning in a crossmodal setup.
@@ 2.4 NLI-based Evaluation
NLI models are trained to determine the relationship between a premise and a hypothesis. Existing approaches either identify a sentence in the reference text as the premise (sentence-level NLI) Nie et al. (2019b); Laban et al. (2022), or measure entailment using the entire reference as the premise Dziri et al. (2022); Honovich et al. (2022). However, a single sentence may lack sufficient content, and using the entire reference is inefficient for long texts Schuster et al. (2022). Context-level NLI addresses this by retrieving relevant sentences to create a short context for NLI Nie et al. (2019a); Schuster et al. (2022); Kamoi et al. (2023), although it still does not operate at a sufficiently detailed or precise level of granularity Nenkova et al. (2007). Here, we propose a novel atomic-level NLI evaluation for the chemical domain to tackle such limitations.
@@ 3.1 Task Definition
Let $(x,y)$ represent a pair of source and target sequences mapped to the $\mathrm{X}$ and $\mathrm{Y}$ spaces, respectively. We cast molecule caption generation (MoCG) as a crossmodal alignment task that operates on offline preference data $\mathcal{D}=\{x^{(i)},y_{w}^{(i)},y_{l}^{(i)}\}_{i=1}^{N}$, where $x$ is the input, and $y_{w}$ and $y_{l}$ are the preferred and dispreferred outputs, respectively, with $N$ being the total number of pairs in $\mathcal{D}$. The goal is to learn an optimal function $f:X\leftrightarrow Y$ via a model $\pi_{\theta}$ parameterised by $\theta$. We coordinate the molecule and caption generation tasks via instruction modelling  [^1] Instructions can be found in Appx. D..
@@ 3.2.1 Universal Models via Model Merging
Let $\tau_{1}$ and $\tau_{2}$ represent task vectors  [^2] A task vector $\tau$ represents the model’s parameters $\Theta^{(t)}$ fine-tuned for task $t$ Ilharco et al. (2022). from pretrained molecule and caption generation models. Our goal is to obtain a multitasking cross-modal model $\Theta^{(merge)}$ without accessing training data by exploring weight-based and subspace-based merging techniques. Fig. 2 illustrates the process.
@@
We experiment with SLERP Goddard et al. (2024), which applies spherical interpolation to fuse model parameters. The goal is to find optimal coefficients $\lambda_{1}$ and $\lambda_{2}$ so that the merged model $\Theta^{(merge)}=\lambda_{1}\tau_{1}+\lambda_{2}\tau_{2}$ retains the capabilities of the independent models. The coefficients are given by $\frac{\sin((1-\lambda)\cdot\rho)}{\sin(\rho)}\quad\text{and}\quad\frac{\sin(\lambda\cdot\rho)}{\sin(\rho)}$, respectively, where $\rho=\arccos\left(\frac{\tau_{1}\cdot\tau_{2}}{|\tau_{1}|\cdot|\tau_{2}|}\right)$ is the angle between the task vectors, and $\lambda$ is the merging coefficient.
@@ 3.2.2 Crossmodal Alignment Fine-tuning
Let $\pi_{ref}$ be the reference policy (i.e., the universal model from model merging), $\pi_{\theta}$ the policy model being trained, parameterised by $\theta$, and $\mathcal{D}=\{x^{(i)},y_{w}^{(i)},y_{l}^{(i)}\}$ the offline preference data. Our goal is to learn effective crossmodals for the MoCG task with minimal training via alignment fine-tuning. Table 1 provides an overview of the optimisation methods employed in this process.
@@
  • CPO Xu et al. (2024) reduces reliance on high-quality data by avoiding suboptimal translations. It modifies Eq. 3 using a uniform reference model, ensuring equal likelihood for all outputs. A behaviour cloning (BC) regulariser is injected to reflect uniform output matching, with an additional SFT term in the final loss (Eq. 4).
@@
  • KTO Ethayarajh et al. (2024) utilises non-paired preference data $\mathcal{D}=\{x^{(i)},y^{(i)},\lambda^{(i)}\}$ where $\lambda$ denotes the desirability of $y$. The loss is computed from the generated output $z$ in relation to a reference $z_{\text{ref}}$ and $\lambda$ (Eq. 5).
@@ 3.3 Atomic-level Cross-NLI Evaluation
Our aim is to develop a method that operates at the right level of granularity, precisely capturing small distinctions and subtle nuances in captions, ensuring reliable evaluation. Atomic-level cross-NLI evaluation uses a LLM and an NLI model to assess relationships between generated and reference captions. The process begins with an LLM Touvron et al. (2023) decomposing a (reference, generated) pair into atomic premises $\{P_{i}\}_{i=1}^{N}$ and hypotheses $\{H_{j}\}_{j=1}^{L}$, where each atomic unit conveys a single piece of information (see Appx. C). An NLI model He et al. (2020) then constructs probabilistic distributions of entailment and contradiction by considering all possible combinations of premises and hypotheses. Finally, pooling operators match atomic hypotheses and premises in terms of both factual correctness, i.e., hallucination, and completeness, i.e., coverage. Fig. 3 illustrates this process.
@@ 4.1 Experimental Setup
We conduct experiments training MeditronChen et al. (2023) on the benchmark L+M-24 Edwards et al. (2024) dataset, using only 10% of the data for training, and evaluate on out-of-distribution data (see Appx. B for details). For alignment fine-tuning, we create synthetic dispreferred outputs generated by MolT5 Edwards et al. (2022). Our training, validation, and test sets contain approximately 12.7k, 3.4k, and 3k samples, respectively. For evaluation, we use established statistical metrics (see Appx. E), along with our atomic-level cross-NLI evaluation method (§ 3.3).
We compare our models thus obtained against established ones trained on extensive data. TxtChem-T5 Christofidellis et al. (2023) is a T5XL multitask model trained on linguistic and molecule modalities across multiple datasets, including CheBI-20, akin to L+M-24. Chem-LLM Zhang et al. (2024), an InternLM2-Base-7B model, is trained on large chemical knowledge databases using DPO, achieving GPT-4-level results. Meditron Chen et al. (2023), a 7B model, is fine-tuned on the entire L+M-24 dataset. Lastly, we fine-tune Meditron with SFT using only 10% of the training data. We leave all the implementation details in Appx. G.

### v3

@@ Abstract
Scientific language models drive research innovation but require extensive fine-tuning on large datasets. This work enhances such models by improving their inference and evaluation capabilities with minimal or no additional training. Focusing on molecule caption generation, we explore post-training synergies between alignment fine-tuning and model merging in a cross-modal setup. We reveal intriguing insights into the behaviour and suitability of such methods while significantly surpassing state-of-the-art models. Moreover, we propose a novel atomic-level evaluation method leveraging off-the-shelf Natural Language Inference (NLI) models for use in the unseen chemical domain. Our experiments demonstrate that our evaluation operates at the right level of granularity, effectively handling multiple content units and subsentence reasoning, while widely adopted NLI methods consistently misalign with assessment criteria.
@@
Dimitris Gkoumas1, Maria Liakata1,2 1Queen Mary University of London, London, UK, 2The Alan Turing Institute, London, UK {d.gkoumas, m.liakata}@qmul.ac.uk
@@ 1 Introduction
AI in Chemistry is essential for developing scalable and cost-effective scientific solutions, such as pioneering drugs Ferguson and Gray (2018), advanced materials Kippelen and Brédas (2009), and improved chemical processes Zhong et al. (2023). The vast search spaces in which these solutions reside make chemical language models crucial for accelerating scientific discovery AI4Science and Quantum (2023); Zhang et al. (2023). Recent trends have led to the use of multimodal models to learn molecular and linguistic representations, either in separate but coordinated spaces Edwards et al. (2021, 2022); Liu et al. (2023a), in a common space Liu et al. (2023b), or through dual approaches Luo et al. (2023); Christofidellis et al. (2023). These models often rely heavily on extensive supervised fine-tuning. However, merely increasing model size and data does not guarantee improvement Tirumala et al. (2022); Xu et al. (2023). Thus, we propose focusing on novel training methods.
[…]
@@
Here we enhance molecule language models using minimal post-training by leveraging synergies between alignment fine-tuning Ouyang et al. (2022) and model merging Yang et al. (2024) in a crossmodal setup. Specifically, we focus on molecule-language translation, using as little as 10% of the training data Edwards et al. (2024). Fig. 1 illustrates our comprehensive post-training solution.
@@
Model merging, a technique for fusing models fine-tuned on different tasks, builds a versatile model without needing the original training data or expensive computation. This method has been quickly adopted in foundation language models Yang et al. (2024). We extend this concept to a crossmodal setting by merging per-task pretrained molecule language models (see Fig. 1), deploying both weight- and subspace-based techniques to obtain universal models (§ 3.2.1).
@@
For fine-tuning alignment, we focus on Reinforcement Learning from Human Feedback (RLHF) Stiennon et al. (2020) to align the universal models. Although alignment has typically been used to calibrate LLM behaviour Askell et al. (2021), we hypothesise that it can also accelerate learning in crossmodal spaces by rewarding preferred over dispreferred outputs, thus improving inference with minimal training data. We focus on optimisation algorithms using closed-form losses on offline preferences, such as Direct Preference Optimisation (DPO) Rafailov et al. (2024), Contrastive Preference Optimisation (CPO) Xu et al. (2024), and Kahneman-Tversky Optimisation (KTO) Ethayarajh et al. (2024). We incorporate golden data as human preferences and dispreferred synthetic outputs generated by proprietary models into the reward signal (see Fig. 1).
@@
We evaluate our models on out-of-distribution data using established statistical-based metrics Sets (2022); Edwards et al. (2022). Additionally, we use Natural Language Inference (NLI) models to assess generated text within the chemical domain. However, off-the-shelf NLI models are suboptimal because: a) they are trained on short texts Williams et al. (2018), while generated outputs may mix overlapping content units Nenkova et al. (2007); b) they struggle with unseen domains McIntosh et al. (2024); and c) they lack subsentence inference, limiting their handling of reordered content (see Fig. 3). Thus, we propose a novel atomic-level cross-NLI approach that addresses these issues. By decomposing reference and generated texts into atomic premises and hypotheses using an LLM, we calculate probability distributions of contradiction and entailment via an NLI model and finally apply row-wise operations to obtain novel hallucination and coverage metrics (§$\S$3.3).
Our findings and contributions are as follows:
@@ 2.1 LLMs for Chemistry
Existing approaches for LLMs in the chemical domain typically rely on costly pretraining with large unimodal datasets for reaction prediction and retrosynthesis Schwaller et al. (2019); Vaucher et al. (2020), or task-specific fine-tuning for language-molecule learning Edwards et al. (2021, 2022, 2024) and molecule editing Liu et al. (2023a); Fang et al. (2023). Other methods focus on multitask learning, which requires resource-intensive pretraining and large multitask datasets Lu and Zhang (2022); Ross et al. (2022); Christofidellis et al. (2023); Zhang et al. (2024). In contrast, we investigate synergies between fine-tuning alignment Gkoumas (2024) and model merging to enhance molecule language models with minimal training.
@@ 2.2 Model Merging
Existing model merging techniques can be broadly categorised into weight-based, subspace-based, and routing-based approaches. Weight-based methods often use optimisation algorithms Yang et al. (2023); Akiba et al. (2024) or geometric interpolations Zhou et al. (2024); Goddard et al. (2024) to determine optimal task vector coefficients. Subspace-based methods involve pruning Yadav et al. (2023); Yu et al. (2024) or masking Wang et al. (2024) to remove insignificant parameters, reducing task interference. Routing-based methods combine models adaptively during inference based on specific input Muqeeth et al. (2023); Tang et al. (2024). We experiment with weight- and subspace-based merging in a crossmodal context.
@@ 2.3 Aligning LLMs
LLM alignment methods can be divided into test-time and fine-tuning approaches. Test-time alignment techniques, such as prompt engineering and guided decoding Khanov et al. (2024); Huang et al. (2024), adjust LLMs without changing their weights, but depend on the original model’s performance. Fine-tuning methods, like RLHF Stiennon et al. (2020); Ouyang et al. (2022), are effective but complex, requiring model retraining and continuous sampling. DPO Rafailov et al. (2024) simplifies RLHF by directly optimizing PPO’s objective, while CPO Xu et al. (2024) improves efficiency by using a uniform reference model. Other methods leverage SFT for optimizing RLHF management and parameter tuning Ethayarajh et al. (2024); Meng et al. (2024). Here, we explore alignment fine-tuning in a crossmodal setup.
@@ 2.4 NLI-based Evaluation
NLI models determine the relationship between a premise and a hypothesis. Existing approaches either identify a sentence in the reference text as the premise (sentence-level NLI)Nie et al. (2019b); Laban et al. (2022), or use the entire reference as the premise Dziri et al. (2022); Honovich et al. (2022), which can be inefficient for long texts Schuster et al. (2022). Context-level NLI addresses this by retrieving relevant sentences to create a short context Nie et al. (2019a); Schuster et al. (2022); Kamoi et al. (2023), but lacks sufficient granularity Nenkova et al. (2007). We propose a novel atomic-level NLI evaluation for the chemical domain to address these limitations.
@@ 3.1 Task Definition
Let $(x,y)$ represent a pair of source and target sequences mapped to the $\mathrm{X}$ and $\mathrm{Y}$ spaces, respectively. We cast molecule caption generation (MoCG) as a crossmodal alignment task that operates on offline preference data $\mathcal{D}=\{x^{(i)},y_{w}^{(i)},y_{l}^{(i)}\}_{i=1}^{N}$, where $x$ is the input, and $y_{w}$ and $y_{l}$ are the preferred and dispreferred outputs, respectively, with $N$ being the total number of pairs in $\mathcal{D}$. The goal is to learn an optimal function $f:X\leftrightarrow Y$ via a model $\pi_{\theta}$ parameterised by $\theta$. We coordinate the molecule and caption generation tasks via instruction modelling  [^1] Instructions can be found in Appx. F..
@@ 3.2.1 Universal Models via Model Merging
Let $\tau_{1}$ and $\tau_{2}$ represent task vectors  [^2] A task vector $\tau$ represents the model’s parameters $\Theta^{(t)}$ fine-tuned for task $t$ Ilharco et al. (2022). from pretrained molecule and caption generation models. Our goal is to obtain a multitasking cross-modal model $\Theta^{(merge)}$ without accessing training data by exploring weight-based and subspace-based merging techniques. Fig. 2 illustrates the process. Specifically, we experiment with model merging approaches that inherently manage conflicts and mitigate modality dominance or instability when integrating modality-specific information using off-the-shelf LLMs, ensuring that neither modality overshadows the other.
@@
We experiment with SLERP Goddard et al. (2024), which applies spherical interpolation to fuse model parameters in a more nuanced approach, blending models in a way that preserves unique characteristics. The goal is to find optimal coefficients $\lambda_{1}$ and $\lambda_{2}$ so that the merged model $\Theta^{(merge)}=\lambda_{1}\tau_{1}+\lambda_{2}\tau_{2}$ retains the capabilities of the independent models. The coefficients are given by $\frac{\sin((1-\lambda_{1})\cdot\rho)}{\sin(\rho)}\quad\text{and}\quad\frac{\sin(\lambda_{2}\cdot\rho)}{\sin(\rho)}$, respectively, where $\rho=\arccos\left(\frac{\tau_{1}\cdot\tau_{2}}{|\tau_{1}|\cdot|\tau_{2}|}\right)$ is the angle between the task vectors, and $\lambda$ is the merging coefficient.
@@ 3.2.2 Crossmodal Alignment Fine-tuning
Let $\pi_{ref}$ be the reference policy (i.e., the universal model from model merging), $\pi_{\theta}$ the policy model being trained, parameterised by $\theta$, and $\mathcal{D}=\{x^{(i)},y_{w}^{(i)},y_{l}^{(i)}\}$ the offline preference data. Our goal is to learn effective crossmodals for the MoCG task with minimal training via alignment fine-tuning. We experiment with different optimizations that differ substantially in how they learn a reward signal, as overviewed in Table 1.
@@
  • CPO Xu et al. (2024) reduces reliance on high-quality data by avoiding suboptimal translations, but not perfect translations in ML tasks. It modifies Eq. 3 using a uniform reference model, ensuring equal likelihood for all outputs. A behaviour cloning (BC) regulariser is injected to reflect uniform output matching, with an additional SFT term in the final loss (Eq. 4).
@@
  • KTO Ethayarajh et al. (2024) utilises non-paired preference data $\mathcal{D}=\{x^{(i)},y^{(i)},\lambda^{(i)}\}$ where $\lambda$ denotes the desirability of $y$. It directly maximizes the utility of generations instead of maximizing the log-likelihood of preferences. The loss is computed from the generated output $z$ in relation to a reference $z_{\text{ref}}$ and $\lambda$ (Eq. 5).
@@ 3.3 Atomic-level Cross-NLI Evaluation
Our aim is to develop a method that operates at the right level of granularity, precisely capturing small distinctions and subtle nuances in captions, ensuring reliable evaluation. Atomic-level cross-NLI evaluation uses an LLM and an NLI model to assess relationships between generated and reference captions. The process begins with an LLM Touvron et al. (2023) decomposing a (reference, generated) pair into atomic premises $\{P_{i}\}_{i=1}^{N}$ and hypotheses $\{H_{j}\}_{j=1}^{L}$, where each atomic unit conveys a single piece of information (see Appx. E). An NLI model He et al. (2020) then constructs probabilistic distributions of entailment and contradiction by considering all possible combinations of premises and hypotheses. Finally, pooling operators match atomic hypotheses and premises in terms of both factual correctness, i.e., hallucination, and completeness, i.e., coverage. Fig. 3 illustrates this process.
@@ 4.1 Experimental Setup
Data:
We conduct experiments training Meditron Chen et al. (2023) on the benchmark L+M-24 Edwards et al. (2024) dataset, using only 10% of the data for training, and evaluate on out-of-distribution data (see Appx. D for details). For alignment fine-tuning, we create synthetic dispreferred outputs generated by MolT5 Edwards et al. (2022). In practice, this involves feeding MolT5 with inputs from the 10% subset of L+M-24 used in our experiments, generating outputs, and then using these outputs as dispreferred samples (see Fig. 1 ). Our training, validation, and test sets contain approximately 12.7k, 3.4k, and 3k samples.
Baselines:
We selected established baselines based on their relevance to our hypotheses, enabling comparison with models trained on fully (i.e., Chem-LLM Zhang et al. (2024)) and partially (i.e., TxtChem-T5 Christofidellis et al. (2023)) out-of-distribution data, as well as in-distribution data (Meditron Chen et al. (2023)). In this context, TxtChem-T5 and Chem-LLM are evaluated in a zero-shot setting. For more details about the baselines, please refer to Appx. G. Lastly, we fine-tune Meditron with SFT using only 10% of the training data. We leave all the implementation details in Appx. J.
Evaluation:
When evaluating the performance of both baselines and our models, we employ established statistical metrics (see Appendix H), in addition to our atomic-level cross-NLI evaluation method (§ 3.3). For our proposed evaluation, we assess the robustness of different NLI methods by measuring the relative entropy of textual entailment between generated outputs from high and low performance models in association with linguistic ones derived by bioinformatic databases curated by humans. Specifically, we compare our atomic-level NLI approach with leading ones, including full NLI, which treats entire premises and hypotheses as single units, and sentence-level NLI Laban et al. (2022), i.e., which evaluates chunks in text.